An object is launched with initial velocity of 22.5 meters per second (m/s) from a 75-meter tall platform. The equation for the object's height h at time t seconds after launched is h(t)=-4.9t^2+22.5t+75, where h is in meters. How long will it take for the object to reach the ground? what is the maximum height reached by the object?
Answer by nerdybill(7384)

An object is launched with initial velocity of 22.5 meters per second (m/s) from a 75-meter tall platform. The equation for the object's height h at time t seconds after launched is h(t)=-4.9t^2+22.5t+75, where h is in meters. How long will it take for the object to reach the ground?
set h(t) to zero and solve for t:
0=-4.9t^2+22.5t+75
applying the quadratic formula, we get:
t = {-2.24, 6.83}
toss out the negative answer leaves:
t = 6.83 seconds
.
what is the maximum height reached by the object?
.
The vertex is the max height. It reaches this at:
t = -b/(2a)
t = -22.5/(2*(-4.9))
t = -22.5/(-9.8)
t = 2.30 seconds
.
h(2.30) = -4.9(2.30)^2+22.5(2.30)+75 = 100.83 meters (max height)